1.25 mol NOCl was placed in a 2.50 L reaction chamber at 427°C. After equilibrium was reached, 1.10 moles of NOCl remained. Calculate the equilibrium constant Kc for the reaction.

2NOCl(g)= 2NO(g) + Cl2(g) 


1
Expert's answer
2021-02-09T03:35:47-0500

Q161077

1.25 mol NOCl was placed in a 2.50 L reaction chamber at 427°C. After equilibrium was reached, 1.10 moles of NOCl remained. Calculate the equilibrium constant Kc for the reaction.

2NOCl(g)= 2NO(g) + Cl2(g) 



Solution:


1.25 mol NOCl was placed in a 2.50 L reaction chamber at 427°C.

So the initial concentration of NOCl is


[NOCl] initial "= \frac{initial \space moles \space of \space NOCl}{volume\space of \space container\space in\space 'L' }" "= \frac{1.25 \space mol \space NOCl}{2.50 \space L}"


"= 0.5 \space mol/L"


When the reaction reached equilibrium, 1.10 moles of NOCl remained in the chamber, so the

concentration of NOCl at equilibrium is


[NOCl]equilibrium "= \frac{equilibrium \space moles \space of \space NOCl}{volume\space of \space container\space in\space 'L' }" "= \frac{1.10 \space mol \space NOCl}{2.50 \space L}"


"= 0.44 \space mol/L"



Next we will draw the ICE table for the given reaction.

I will show the image of the ICE table I have drawn.





So we have


[NOCl] equilibrium = 0.5 – x = 0.5 – 2 * 0.03  = 0.5 – 0.06 = 0.44 mol/L


[NO] equilibrium  =  2x = 2 * 0.03   = 0.06 mol/L


[Cl2] equilibrium = x = 0.03 mol/L 


Now consider the reaction


2NOCl (g) <==> 2NO(g) + Cl2(g)

equilibrium constant equation for this reaction is written as


"Kc = \frac{[NO]^2*[[Cl_2]}{[NOCl]^2}" ;


plug the equilibrium concentration in this equation we have


"Kc = \frac{[0.06mol/L]^2*0.03mol/L}{[0.44mol/L]^2}" "= \frac{0.0036*0.03 }{0.1936} = \frac{0.000108}{0.1936} = 0.0005579 mol/L"



which can also be written as Kc = 5.58 * 10-4 mol/L ;


So the equilibrium constant, Kc for the given reaction is 5.58 * 10-4 mol/L